OpenStudy (anonymous):

3a - 3b = -15 -3a - 3b = -3

OpenStudy (adamaero):

What method do you want to use? Substitution or elimination?

OpenStudy (adamaero):

I guess I should say which*

OpenStudy (anonymous):

substitution

OpenStudy (adamaero):

1st step: Solve one equation for one variable.

OpenStudy (anonymous):

3a - 3b = -15 -3a - 3b = -3 ________________ -6b = -18 b= 18/6, b=3 put this value in any eq i will use the first 3a - 3(3) = - 15 3a=-15+9, 3a=-6 a=-2 this method would be more easier